# MICL: Improving In-Context Learning through Multiple-Label Words in Demonstration

### v1

@@ 4.5 Effectiveness of MICL Ordering
Table 5: Effectiveness of MICL’s Order: Comparison of our ordering method with random orders in multi-class datasets and the flipped order in binary datasets. ‘Permutation’ presents the best result among the evaluated permutations for each dataset.
@@
As shown in Table 5, MICL often outperforms or matches the best results among compared permutations. It achieves an average accuracy improvement of 0.39% in Llama2-7b and 5.86% in GPT2-xl. Despite a notable performance drop for IMDB (GPT2-xl), the binary task nature can mitigate this by enumerating all orders. MICL excels in multi-class tasks, matching or exceeding top permutation results. This highlights the effectiveness of MICL’s demonstration order, particularly in multi-class tasks with thousands of possible orderings ($7!$), which would take months to process enumerating in LLMs, whereas MICL achieves comparable performance within hours or minutes.
4.6 Effectiveness of Label Balance in MICL
SST2 CR IMDB TREC AMAN ISEAR AGNews Avg.
Llama2-7b
MICL 95.39 94.41 95.40 78.40 59.90 72.49 84.11 82.87
MICL
unbalanced
91.98 93.35 93.60 68.00 57.39 70.70 81.49 79.50
+Demo-MLabels_CN 95.97 95.15 95.60 79.80 65.16 73.55 86.55 84.54
+Demo-MLabels_CN
unbalanced
92.42 93.62 94.40 69.60 62.16 71.36 85.87 81.35
+Demo-MLabels_LW 95.97 95.15 95.60 80.60 69.40 73.09 86.58 85.20
+Demo-MLabels_LW
unbalanced
92.42 93.62 94.40 81.80 62.91 71.36 86.36 83.27
GPT2-xl
MICL 85.17 64.89 71.50 61.60 47.87 48.64 78.92 66.71
MICL
unbalanced
51.02 73.67 67.80 53.60 41.60 54.62 34.45 53.82
+Demo-MLabels_CN 91.65 65.96 73.40 70.40 49.62 59.73 79.08 70.55
+Demo-MLabels_CN
unbalanced
52.85 74.47 68.20 55.40 45.86 56.75 43.75 56.75
+Demo-MLabels_LW 91.65 65.96 73.40 70.40 49.87 58.74 79.49 70.50
+Demo-MLabels_LW
unbalanced
52.85 74.47 68.40 59.60 45.61 56.42 43.75 57.30
Table 6: Accuracy performance (%) of MICL, ‘+Demo-MLabels_CN’, ‘+Demo-MLabels_LW’ under label-balanced and label-unbalanced demonstrations.
The impact of label-unbalanced demonstrations varies between the two models. For Llama2-7b, a large-size language model, the negative effects of missing label information are less marked, with an average accuracy decrease of 3.37%. The most significant accuracy drops in TREC by 10.40%. The leverage of multiple-label words mitigated the performance loss, reducing the average accuracy decline to 1.93% in ‘+Demo-MLabels_LW’ compared to a label-balanced setting. In contrast, GPT2-xl, a smaller model with 1.5 billion parameters, experienced a marked performance decline under unbalanced conditions, averaging a 12.89% decrease in accuracy. Specifically, SST2 and AGNews experienced over 34% declines. Interestingly, CR and ISEAR demonstrated improved performance despite the unbalanced labels. The incorporation of multiple-label words consistently enhanced performance across all datasets in the unbalanced setting, affirming their utility in ICL.
4.7 Label Word Filtering Results
Table 7 lists the number of filtered words (-) for each dataset under Llama2-7b and GPT2-xl during distribution separability filtering ($S_{1}$) and Point-Biserial testing filtering ($S_{r}$).
Model Filtered Verbalizer SST2 CR IMDB TREC AMAN ISEAR AGNews
Llama2-7b $S_{1}$315 -228 -231 -31 -75 -78 -1163
$S_{r}$1 -9 -9 -1 -3 0 0
GPT2-xl $S_{1}$393 -368 -383 -40 -62 -118 -645
$S_{r}$4 -7 -6 -2 -6 0 -3
The large number of words filtered by distribution separability indicates that although many words match the task topic definition at the linguistic level, they are not suitable as label words at the LLM level. This suggests that simple label-based voting, commonly used in many prompt-based methods, might harm LLM’s in-context learning, as the candidate words do not align with the task based on the LLM’s understanding. The small number of words filtered by Point-Biserial testing indicates the high quality of the proposed distribution separability filtering, as the remaining words show significant separation between the true class and false classes based on logit vectors.
@@ 5 Conclusion
This paper introduces MICL, a novel approach to organizing demonstrations with multiple-label words inserted in in-context learning (ICL). By utilizing a variety of label words and analyzing their distribution within large language models (LLMs), we enhance ICL understanding by providing diverse label information. We develop a structured method for selecting and ordering sample-multiple-label pairs via LLM’s feedback over label words. Extensive experimental results show that our method of multiple-label word insertion significantly improves ICL classification performance, yielding superior results.
@@ 6 Limitations
In this paper, we enhance in-context learning performance by incorporating additional label-related words. Although related label words for various tasks have been extracted and collected, new datasets may still lack appropriate label words. However, powerful search tools such as WordNet Pedersen et al. (2004), ConceptNet Speer et al. (2017), and open-source vocabularies can mitigate this issue. Our method designs a filtering approach that refines the quality of label words based on these search results.
@@ Acknowledgements
Appendix A Appendix A
A.1 Label Effectiveness in ICL
This study establishes four distinct label sets for each dataset, utilizing identical samples to form the sample-label pairs in 1-shot ICL, as detailed in Table 8. The reported results represent the average accuracy obtained from five repeated experiments, conducted with seeds 42, 43, 44, 45, and 46 in sample selection during the 1-shot demonstrations. In Llama2-7b, the maximum 49.20% accuracy difference (TREC), and the highest 11.08% (SST2) standard deviation are observed. We also evaluate the label effectiveness in GPT2-xl under the same experimental conditions, with results shown in Fig. 4. The results are similar to those in Llama2-7b, with a maximum accuracy difference of 48.20% (TREC) and a maximum standard deviation of 10.52% (SST2). These findings indicate that label selection contributes to both ICL accuracy and robustness.
@@ A.1 Label Effectiveness in ICL
3: ‘ human’, 4: ‘ location’,5: ‘location’
@@
Figure 4: Label effectiveness in ICL (GPT2-xl)
A.2 LLM’s Output Separability Over Label Words in Zero-Shot Learning
This study evaluates the logits value separability for negative and positive samples of label words ‘bad’ and ‘pessimistic’ compared to the class name ‘negative’, and for label words ‘good’ and ‘happy’ compared to the class name ‘positive’ in SST-2 under zero-shot learning. Except for the word bad’, shown in Fig. 1(b), the logit distribution figures for the remaining words are listed in Fig 5.
@@ A.2 LLM’s Output Separability Over Label Words in Zero-Shot Learning
Figure 5: Label words logit separability over samples.
In Fig. 1(b) and Fig. 5, the first 100 samples are negative, while samples 101-200 are positive samples. Compared to the class names, the logits of negative label words across negative samples are higher than those for ‘negative’, while the logits for the same label words across positive samples are lower than ‘negative’. Similarly, the logits for positive label words are higher across positive samples and lower across negative samples than ‘positive’, indicating better logit separability for these label words compared to their respective class names. This demonstrates superior logit separability for certain label words compared to class names. Since logit values are crucial for class prediction, this enhanced separability can significantly improve classification performance.
A.3 Multiple label words effectiveness in ICL
@@
This study evaluates the performance of demonstrations using different numbers of label words. The multiple-label words combine the class name with related label words, connected by spaces, such as "negative bad" and "positive good" in SST2 and CR. We also assess the label effectiveness in GPT2-xl under the same experimental conditions, with the results shown in Fig. 6. These findings indicate the potential of using multiple-label words in demonstrations to enhance ICL.
Figure 6: Multiple label words effectiveness in ICL (GPT2-xl)
Appendix B Appendix B
@@
The label word sets for SST2, CR, IMDB, and AGNews are derived from Hu et al. (2022), while those for TREC, AMAN, and ISEAR are sourced from Zhu et al. (2024).
@@
Appendix C Appendix C
@@
Baseline Model Experimental Settings The detailed information on the baseline models and the corresponding experimental settings for few-shot learning experiments is provided below.
@@
For all experiments conducted with Llama2-7b, the model is configured to operate under a 4-bit setting.
Appendix D Appendix D
The Number of Inserted Labels Settings The number of label words (N) used in Table 1 on the baseline models are summarized in Table 10.
@@ Appendix D Appendix D
MICL 2 2 2 2 2 7 2
Table 10: The number of label words (N) inserted in demonstration in the baseline models and MICL (ours) under each dataset in 1-shot ICL.
@@
Figure 7: Validation and test performance under different label word quantity (N) in sample-multiple-label pairs for CR. IMDB, ISEAR and AGNews.
The number of label words (N) used in Table 3 on the baseline models are summarized in Table 11.
@@
Table 11: The number of label words (N) inserted in demonstration in SelfICL and MICL (ours) under SST2 and AMAN in 5-shot ICL.

### v4

@@ 4.4.1 Effectiveness of LICL Ordering
Table 4: Accuracy (%) comparison of our ordering method with random orders in multi-class datasets and the flipped order in binary datasets. ‘Permutation’ presents the best result for each dataset.
@@
As shown in Table 4, LICL often outperforms or matches the best results among compared permutations. It achieves an average accuracy improvement of 0.39% in LLaMA2-7b and 5.86% in GPT2-xl. Despite a notable performance drop for IMDB (GPT2-xl), the binary task nature can mitigate this by enumerating all orders. LICL excels in multi-class tasks, matching or exceeding top permutation results. This highlights the effectiveness of LICL’s demonstration order, particularly in multi-class tasks with thousands of possible orderings ($7!$), which would take months to process enumerating in LLMs, whereas LICL achieves comparable performance within hours or minutes.
4.4.2 Effectiveness of Label Organization
In Table 1, incorporating multiple class-related words significantly boosts ICL performance across all models. We assess our label selection and ordering method by comparing it against a random organization of class-related words. In Table 5, the observed decreased accuracy in ‘+MLabels-Random’ underscores the superiority of our systematic approach.
LICL LLaMA2-7b SST2 CR IMDB TREC AMAN ISEAR AGNews
+MLabels-Random 94.84 93.62 94.90 79.40 57.64 70.56 80.49
+MLabels_LW 95.97 95.15 95.60 80.60 69.40 73.09 86.58
Table 5: Accuracy (%) comparison of our organization method with random organization in $P_{r}$ (same N as in Table 1).
4.5 Scalability to Larger LLMs
Table 6 presents the accuracy results of LICL implemented on LLaMA3-8b and LLaMA2-13b to further validate the effectiveness and applicability of our method on larger language models. The insertion of multiple class-related words led to an increase in accuracy. LICL achieved a 3.42% accuracy improvement on LLaMA3-8b and a 2.32% improvement on LLaMA2-13b. Specifically, for the IMDB dataset on LLaMA2-13b, despite significant sample information loss compared to other methods, our approach maintained strong performance. Details of the evaluation of initial class-related word updates can be found in Appx.G.2.
SST2 CR IMDB TREC AMAN ISEAR AGNews Avgs.
LLaMA3-8b
LICL 95.66 94.29 95.50 79.00 61.15 70.83 84.67 83.01
+MLabels_CN 95.77 95.82 95.90 80.40 65.89 73.89 87.72 85.06
+MLabels_LW 95.77 95.82 95.90 86.00 69.57 74.29 87.72 86.44
LLaMA2-13b
LICL 96.38 94.62 95.70* 82.20 60.15 73.16 85.12 83.90
+MLabels_CN 96.87 95.88 95.90* 84.00 67.64 74.37 87.79 86.06
+MLabels_LW 96.87 95.88 95.90* 84.40 67.89 74.82 87.79 86.22
Table 6: Accuracy performance (%) of LICL on LLaMA2-13b and LLaMA3-8b. *: For the IMDB dataset on LLaMA2-13b, our GPU could only accommodate 20% of each demonstration sample length.
4.6 Inserted Words’ General Applicability
To address potential concerns about overfitting when selecting multiple class-related words from a single dataset, we implemented cross-dataset evaluations. We evaluated the efficacy of using two class-related words (N in most binary tasks in our study) selected from the SST2 dataset across the CR and IMDB datasets, as detailed in Table 7.
CR IMDB CR IMDB
LLaMA2-7b GPT2-xl
LICL 94.41 95.40 64.89 71.50
+MLabels (from same dataset) 95.15 95.60 65.96 73.40
+MLabels (from SST2) 95.15 95.40 65.04 72.86
Table 7: Cross-dataset accuracy (%) evaluation on inserted multiple class-related words.
In LLaMA2-7b, performance improvements on the CR may be attributed to similar sub-tokens found in both SST2 and CR (e.g., " unhealthy", " unfair", " good"). No performance decrease is observed in IMDB. In GPT2-xl, despite initial differences in the positive sentiment words across datasets, the class-related words from SST2 still enhanced performance in both CR and IMDB. These results suggest that the selected class-related words exhibit a level of generality, enhancing ICL performance across various datasets with similar labels.
4.7 Further Analyses
More in-depth analyses are in Appx.H, including pool filtering effectiveness and its ablation study (H.1), label bias evaluation (H.2), the impact of class-related words on predictions whether or not they appear in demonstrations (H.3), the effects of label-balanced demonstration in LICL (H.4), inference cost (H.5), and sample-label distribution visualizations (H.6).
@@ 5 Conclusion
In this paper, we introduce logit separability, a criterion for measuring the class-related clarity of samples and labels based on their logit value distributions, assessing whether a sample is easily predictable by its true label and whether a label performs well across multiple samples of the same class. To improve ICL with clearer class-related guidance, we propose LICL, a logit separability-based method for organizing demonstrations. Besides, to enrich ICL with broader label information, we present a novel label utilization method by forming sample-multiple-label pairs through the insertion of multiple class-related words, based on the logit distribution of selected samples. Experimental results confirm that this combined approach significantly enhances ICL classification performance, yielding superior outcomes.
@@ 6 Limitations
In this paper, we enhance in-context learning performance by incorporating additional class-related words. Although related class-related words for various tasks have been extracted and collected, new datasets may still lack appropriate class-related words. However, powerful search tools such as WordNet Pedersen et al. (2004), ConceptNet Speer et al. (2017), and open-source vocabularies can mitigate this issue. Our method designs a filtering approach that refines the quality of class-related words based on these search results.
@@ Acknowledgements
Appendix A Exploring Logit Separability and Class-Related Word Effectiveness
A.1 Logit Separability Over Class-Related Words in Zero-Shot Learning
This study evaluates the logit value separability for negative and positive samples of class-related words ‘bad’ and ‘pessimistic’ compared to the class name ‘negative’, and for class-related words ‘good’ and ‘happy’ compared to the class name ‘positive’ in SST-2 under zero-shot learning. Except for the word bad’, shown in Fig.1(b), the logit distribution figures for the remaining words are listed in Fig 4.
@@ A.1 Logit Separability Over Class-Related Words in Zero-Shot Learning
3: ‘ human’, 4: ‘ location’,5: ‘ number’
@@
In Fig.1(b) and Fig.4, the first 100 samples are negative, while samples 101-200 are positive samples. Compared to the class names, the logit of negative class-related words across negative samples is higher than those for ‘negative’, while the logit for the same class-related words across positive samples is lower than ‘negative’. Similarly, the logit for positive class-related words is higher across positive samples and lower across negative samples than ‘positive’, indicating better logit separability for these class-related words compared to their respective class names. This demonstrates superior logit separability for certain class-related words compared to class names. Since logit values are crucial for class prediction, this enhanced separability can significantly improve classification performance.
@@
Figure 4: class-related words logit separability over samples.
A.2 Label Effectiveness in ICL
A.2.1 Single Class-Related Word
This study establishes four distinct label sets for each dataset, utilizing identical samples to form the sample-label pairs in 1-shot ICL, as detailed in Table 8. The reported results represent the average accuracy obtained from five repeated experiments, conducted with seeds 42, 43, 44, 45, and 46 in sample selection during the 1-shot demonstrations. As shown in Fig.5, in LLaMA2-7b, the maximum 49.20% accuracy difference (TREC), and the highest 11.08% (SST2) standard deviation are observed. We also evaluate the label effectiveness in GPT2-xl under the same experimental conditions, with results shown in Fig.6. The results are similar to those in LLaMA2-7b, with a maximum accuracy difference of 48.20% (TREC) and a maximum standard deviation of 10.52% (SST2). These findings indicate that label selection contributes to both ICL accuracy and robustness.
Figure 5: Label effectiveness in ICL (GPT2-xl)
Figure 6: Label effectiveness in ICL (GPT2-xl)
A.2.2 Multiple class-related words
@@
This study evaluates the performance of demonstrations using different numbers of class-related words. The multiple class-related words combine the class name with related class-related words, connected by spaces, such as "negative bad" and "positive good" in SST2 and CR. We also assess the label effectiveness in GPT2-xl under the same experimental conditions, with the results shown in Fig.7. These findings indicate the potential of using multiple class-related words in demonstrations to enhance ICL.
Figure 7: Multiple class-related words effectiveness in ICL (GPT2-xl)
Appendix B Class-Related Word Pool Filtering
The goal of class-related word pool filtering is to enhance the logit separability of the words within the pool, guided by logit feedback from LLM across the dataset. Ideally, each word should exhibit a clearly dominant logit value in samples with the corresponding label, compared to those with different labels, to prevent the introduction of incorrect class-related features in the ICL demonstration.
To achieve this, we first evaluate the separability of each class-related word based on its logit value across the training set $D$ in a zero-shot setting, where input samples are formatted using a template without labels. For a word $w\in P$ with label $l\in L$, we exclude it if its average logit $b_{l}$ for training samples labeled with $l$ is not the highest among all training samples. This step ensures that the class-related word provides accurate logit feedback in the corresponding training samples compared to those with different label expressions. To further ensure the separability of the feedback, avoiding situations where, for example, a class-related word has an average logit of 7.1 in correct training samples but a close value of 7.0 in semantically unrelated samples, we apply point-biserial testing to test the correlation $r$ between the word’s logit values in samples with the same label versus those with different labels as defined in Eq.4:
$r=\frac{V_{0}-V_{1}}{\sigma}\times\sqrt{\frac{n_{0}\times n_{1}}{(n_{0}+n_{1})^{2}}}$ (4)
where $V_{0}$ denotes the mean logit values of the class-related word for training samples with the same label, $V_{1}$ denotes the mean logit values for samples with different labels, $\sigma$ represents the standard deviation of the logit values, and $n_{0}$ and $n_{1}$ correspond to the number of samples with the same and different labels, respectively.
For a given class-related word, a positive correlation $r>0$ indicates that its average feedback across training samples within its own class significantly differs from that of samples in other classes, warranting its retention. Otherwise, the word will be removed from the pool. This process ensures that the remaining words in the refined pool $P_{r}$ are not only logit separable but also exhibit clear class-related information distinctions.
Appendix C Prompt Examples
The prompts used in this paper are exampled below:
Zero-Shot Learning:
Review: they ’re easy to use
Sentiment:
One-Shot Learning:
Review: norton support is completely
pathetic
Sentiment: negative
Review: overall , i am very pleased
with it
Sentiment: positive
Review: they ’re easy to use
Sentiment:
One-Shot Multiple Class-Related Word Insertion Learning (result from CR in LlaMA2-7b):
Review: it does not only have
difficulty playing jpegs , it even
has trouble ...
Sentiment: negative unhealthy unjust
Review: about the product the zen
micro is a sleek , stylish ...
Sentiment: positive good favorable
Review: they ’re easy to use
Sentiment:
Appendix D Statistic Information of Datasets and Templates
@@
The class-related word pools for SST2, CR, IMDB, and AGNews are derived from Hu et al. (2022), while those for TREC, AMAN, and ISEAR are sourced from Zhu et al. (2024).
@@
Appendix E Experimental Implementation Details
@@
Baseline Model Experimental Settings The detailed information on the baseline models and the corresponding experimental settings for few-shot learning experiments is provided below. For all experiments conducted with LLaMA2-7b, LLaMA3-13b, and LLaMA3-8b, the model is configured to operate under a 4-bit setting.
@@
Table 10 indicates whether each baseline and our proposed method incorporate functions for the selection and ordering of samples and labels, demonstrating that our one-stop method encompasses all components of demonstration organization.
Sample Label
Selection Ordering Selection Ordering
Vanilla X X X -
TopK X -
SelfICL X -
DataICL X X -
LICL(ours)
Table 10: Assessment of Demonstration Organization Methods in the Selection and Ordering of Samples and Labels. A ‘✓’ indicates that the method incorporates this feature, an ‘X’ signifies its absence, and a ‘-’ indicates that the feature is not applicable.
Appendix F The Number of Inserted Labels Settings
The number of class-related words (N) used in Table 1 on the baseline models are summarized in Table 11.
@@ Appendix F The Number of Inserted Labels Settings
LICL 2 2 2 2 2 7 2
Table 11: The number of class-related words (N) inserted in the demonstration in the baseline models and LICL (ours) under each dataset in 1-shot ICL.
@@
Figure 8: Validation and test performance under different class-related word quantity (N) in sample-multiple-label pairs for CR. IMDB, ISEAR and AGNews.
The number of class-related words (N) used in Table 3 on the baseline models are summarized in Table 12.
@@
Table 12: The number of class-related words (N) inserted in the demonstration in SelfICL and LICL (ours) under SST2 and AMAN in 5-shot ICL.